what is 6% of 120
what % of 8 is 150

when question like these are aske what is the key to knowing when to / or multiply

Answer by Nate(3500)

what is 6% of 120
120*0.06 = 7.2
what % of 8 is 150
8x = 150
x = 150/8 = 18.75 or %1875
You just need to read carefully to understand what you have to do with the number you are trying to discover.